OpenStudy (friendzone):

A method of solving a system of equations in which one variable is replaced by an expression using the other variable as a representation for that variable.

OpenStudy (anonymous):

what's your question?

OpenStudy (michele_laino):

I think you are talking about the substitution method

OpenStudy (friendzone):

its a term

OpenStudy (friendzone):

yeah lol no one ever understands it

OpenStudy (anonymous):

m+n =5 m-n=3 then m=3+n put in first (3+n) +n =5 and solve